a particle moves a distance x in time t according to equation x=(t+5)^-1,the acceleration of the particle is proportional to 
  1. velocity^3/2
  2. distance^2
  3. distance^-2
  4. velocity^2/3

Grade:12th pass

1 Answers

Vikas TU
7 years ago
accleration = duble differentiation of displacemnt w.r.t time.
Hence,
dx/dt = -(t+5)^-2  = v
d2x/dt^2  = 2(t+5)-3 = a
subsitute v in a and get the answer.